# Text-Driven Traffic Anomaly Detection with Temporal High-Frequency Modeling in Driving Videos

### v1

@@ Methods
IV-D Quantitative Results
IV-D1 Overall results
IV-D2 Per-class results
IV-D3 Generalization performance
IV-E Qualitative Results
In this subsection, we present visual examples in Fig. 4 to further illustrate the detection capability of our TTHF framework across various types of traffic anomalies. Specifically, we show five representative traffic anomalies from top to bottom as examples: a) The other vehicle collides with another vehicle that turns into or crosses a road. b) The ego-vehicle collides with another oncoming vehicle. c) The ego-vehicle collides with another vehicle moving laterally in the same direction. d) The ego-vehicle collides with another vehicle waiting. e) The ego-vehicle is out-of-control and leaving the roadway to the left. From the above visualization results of different types of traffic anomalies, we can summarize as follows. Overall, our TTHF exhibits superior detection performance on various types of traffic anomalies. Secondly, while the most intuitive classify-based approach (It has the same network architecture as the visual encoder of TTHF, but directly classifies the visual representation, denoted as Classifier in Fig. 4) also follows a single-stage paradigm, our proposed text-driven TAD approach offers a more comprehensive representation in high-dimensional space than orthogonal one-hot vectors. Consequently, both our proposed TTHF and its variants outperform the Classifier. Third, incorporating AAFM allows our method to better perceive different types of traffic anomalies, as evident in Fig. 4 when comparing the Base and AAFM variants across various traffic anomalies. Finally, capturing dynamic changes in driving scenarios significantly enhances traffic anomaly detection. This highlights the effectiveness of our approach in characterizing dynamic changes in driving scenarios by modeling high frequency in the temporal domain.
IV-F Ablation Investigation
IV-F1 Variants of our architecture
@@ IV-F2 Analysis of the AAFM
Arch. Visual
Textual
AUC (%)
TTHF RN-50
Text-Transformer
84.7
TTHF RN-50x64
Text-Transformer
84.8
TTHF ViT-B-32
Text-Transformer
84.0
TTHF ViT-L-14
Text-Transformer
85.0

### v2

@@ IV-C Competitors
Arch. Visual Textual AUC (%)
TTHF RN-50 Text-Transformer 84.7
TTHF RN-50x64 Text-Transformer 84.8
TTHF ViT-B-32 Text-Transformer 84.0
TTHF ViT-L-14 Text-Transformer 85.0
IV-D Quantitative Results
IV-D1 Overall results
Figure 6: Visualization of the weights used for computing soft text representations. We present three illustrative cases, each involving video frames captured at different times. These frames are accompanied by the corresponding weight values used in the computation of soft text representations. Notably, we employ a blue-to-red color scale, where increasing redness signifies higher weights. Additionally, we label the ground-truth fine-grained text representations (denoted as T_i) associated with specific frames. Among them, T_1 corresponds to the text “The ego vehicle collision with another vehicle” (as described in Table I), T_4 corresponds to the text “The non-ego vehicle collision with another vehicle”, T_7 corresponds to the text “The ego vehicle out-of-control and leaving the roadway”, and T_11 corresponds to the text “The vehicle is running normally on the road”.
IV-D2 Per-class results
IV-D3 Generalization performance
IV-E Qualitative Results
In this subsection, we visualize some examples to further illustrate the detection capability of our TTHF across various types of traffic anomalies and the feasibility of soft text representation in our framework.
IV-E1 Visualization of various types of traffic anomalies
As presented in Fig. 5, we show five representative traffic anomalies from top to bottom as examples: a) The other vehicle collides with another vehicle that turns into or crosses a road. b) The ego-vehicle collides with another oncoming vehicle. c) The ego-vehicle collides with another vehicle moving laterally in the same direction. d) The ego-vehicle collides with another vehicle waiting. e) The ego-vehicle is out-of-control and leaving the roadway to the left. From the above visualization results of different types of traffic anomalies, we can summarize as follows. Overall, our TTHF exhibits superior detection performance on various types of traffic anomalies. Secondly, while the most intuitive classify-based approach (It has the same network architecture as the visual encoder of TTHF, but directly classifies the visual representation, denoted as Classifier in Fig. 5) also follows a single-stage paradigm, our proposed text-driven TAD approach offers a more comprehensive representation in high-dimensional space than orthogonal one-hot vectors. Consequently, both our proposed TTHF and its variants outperform the Classifier. Third, incorporating AAFM allows our method to better perceive different types of traffic anomalies, as evident in Fig. 5 when comparing the Base and AAFM variants across various traffic anomalies. Finally, capturing dynamic changes in driving scenarios significantly enhances traffic anomaly detection. This highlights the effectiveness of our approach in characterizing dynamic changes in driving scenarios by modeling high frequency in the temporal domain.
IV-E2 Visualization of the weights used for soft text representation
We further investigate the feasibility of soft text representations. Specifically, as shown in Fig. 6, we use three cases from the test set as examples. For video frames captured at different moments in driving videos, we visualize the weights employed to compute the soft text representation and compare it with the real fine-grained text representation. From the visualization results, we observe that the text representation associated with the maximum weight (indicated by the darkest red) consistently aligns with the real fine-grained text representation. The above results indicate that the way we calculate the soft text representation is effective and can well reflect the real anomaly category.
IV-F Ablation Investigation
IV-F1 Variants of our architecture